# Role of activity and dissipation in achieving precise beating in cilia: Insights from the rower model

### v1

@@ Abstract
Cilia and flagella are micron-sized filaments that actively beat with remarkable precision in a viscous medium, driving microorganism movement and efficient flow. We study the rower model to uncover how cilia activity and dissipation enable this precise motion. In this model, cilia motion is represented by a micro-bead’s Brownian movement between two distant harmonic potentials. At specific locations, energy pumps trigger potential switches, capturing cilia activity and generating oscillations. We quantify precision of oscillation using a quality factor, identifying its scaling with activity and oscillation amplitude, finding precision maximization at an optimal amplitude. The data collapse is not accurate for noisy oscillations. An exact analytic expression for the precision quality factor, based on first passage time fluctuations, and derived in the small noise approximation, explains its optimality and scaling. Energy budget analysis shows the quality factor’s consistency with the thermodynamic uncertainty relation.
@@ I Introduction
Motile cilia and flagella, hairlike appendages found on many unicellular organisms and epithelial tissues, actively beat to drive movement [1, 2, 3, 4]. Their structure consists of nine outer microtubule doublets and two central doublets [5], with sliding movement powered by dynein motors [6, 7]. This internal force causes microtubules to move along the axoneme [5, 8, 9, 10, 11, 12], generating complex oscillations in the surrounding fluid medium. These movements, often synchronized into metachronal waves, are essential for processes like mucociliary clearance [13, 2, 14] and the locomotion of microorganisms such as Paramecium, Opalina, and Volvox [15, 16, 17].
Cilia and flagella exhibit both coordinated movement in large groups and remarkable precision in their individual oscillations, which is surprising given the noisy environments in which they operate. The precision of these oscillations is quantified using a quality factor $Q$, a dimensionless ratio of coherence time to oscillation period [18, 19, 20, 21]. Values of $Q$ typically range from $\sim 30-100$ [22, 23, 24, 25, 26, 27, 28], indicating high precision. In Chlamydomonas reinhardtii, flagellar precision increases with the number of motor proteins. Experimental findings are summarized in a recent study [28].
@@
We investigate how the interplay between cilia activity and dissipation due to the surrounding viscous medium influences precision of oscillation, a topic of recent interest [27, 28]. To explore this, we first analyze the rower model numerically and then analytically, quantifying precision using two related definitions of quality factor and linking it to fluctuations in the first passage time (FPT). Our results show a clear data collapse and scaling of the quality factor, revealing a competition between activity and effective temperature. However, increasing noise strength causes deviations from the nice data collapse. Precision varies non-monotonically with oscillation amplitude, peaking at an optimal value. Applying a small noise approximation (SNA), we derive a closed-form expression for the FPT fluctuation and hence the quality factor, offering key insights into the role of activity and dissipation in precise beating. This expression explains the observed data collapse and scaling, and captures the precision peak at an optimal oscillation amplitude. As we show, the energy budget for achieving a desired precision follows the thermodynamic uncertainty relation.
@@ II The Rower Model
Here, $k$ is the stiffness, and $\mu$ is the separation between the minima of the two potentials, a natural length scale of the system. The motion is confined to the range ±$\pm\mathcal{A}$, where $2\mathcal{A}<\mu$. The bead undergoes Brownian motion only in the downhill regions of the potentials, and switching between the potentials at the terminal positions drives the bead in the opposite direction, creating sustained oscillations. Thermal fluctuations and other stochastic processes introduce noise into the oscillations. The over-damped motion of the bead is described by:
@@
where $\gamma=6\pi\eta a$ is the coefficient of drag force on the bead with radius $a$ due to the fluid with viscosity $\eta$, and $\xi(t)$ the white noise at time $t$. The noise satisfies $\langle\xi(t)\rangle=0$ and $\langle\xi(t)\xi(t^{\prime})\rangle=2D\delta(t-t^{\prime})$, where $D$ denotes a translational diffusivity. Within this model, the diffusivity and drag coefficient can be used to define an effective temperature
@@
which, for active cilia, arises out of both thermal fluctuations and chemical processes involved in active energy pumping [72, 73, 26, 74, 75]. We investigate temporal coherence in oscillations of this system. When the bead reaches a terminal position, the potential switches, injecting energy $\mu k\mathcal{A}$, which we refer to as the bead’s activity.
Figure 1: Oscillations in the rower model for cilia. (A) A schematic of the realistic beating of a cilium is depicted. (B) It illustrates the oscillation mechanism in the rower model. A micron-sized bead in a fluid medium moves in the downhills of two harmonic potentials (represented by $\sigma=\pm 1$) with stiffness $k$ and separated by a distance $\mu$. The motion is restricted between $-\mathcal{A}$ to $\mathcal{A}$ with $2\mathcal{A}<\mu$. Once the bead reaches a terminal position (±$\pm\mathcal{A}$), switching between the potential happens, implying the pumping of energy $\mu k\mathcal{A}$. (C) and (E) Typical oscillating trajectories for high noise ($D=0.02\mu m^{2}/s$) and low noise ($D=0.005\mu m^{2}/s$) strengths are shown. (D) and (F) Auto-correlation functions for trajectories for subplots (C) and (E) are shown. Other parameters used: $\mu=1\mu m$, $\mathcal{A}=0.25\mu m$, and $k=1.5pN/\mu m$.
Within a single potential well, the bead relaxes to equilibrium with a relaxation time $\tau_{d}=\gamma/k$, a natural time scale, in the absence of switching. The switching causes sustained oscillations of the cilium between $-\mathcal{A}$ and $\mathcal{A}$ in this overdamped dynamics. The average period $T_{0}$ of the oscillations is obtained by solving the mean dynamics for each $\sigma$ separately, and is given by:
@@ III Results
In this section we present definitions of quality factor describing the precision of oscillation and analyze the results, first numerically and then analytically within SNA.
@@ III.2 Precision quality factors of oscillations
Figure 2: The precision quality factor results for various activity and dissipation parameters are shown. Different symbols represent simulation results for different conditions. (A) $Q$ vs. $\mathcal{A}$ for varying $\mu$ ($k=4$ pN/$\mu$m and $D=0.02$). (B) $Q$ vs. $\mathcal{A}$ for varying $k$ ($\mu=2\mu$m and $D=0.02$). (C) $Q$ vs. $\mathcal{A}$ for different noise strengths $D$ ($\mu=2\mu$m, $k=4$ pN/$\mu$m). (D) $Q$ vs. $\mathcal{A}$ for varying $\eta$ under thermal noise ($\mu=3\mu$m, $k=2$ pN/$\mu$m). (E) Scaled $Q$ vs. $\mathcal{A}_{s}=2\mathcal{A}/\mu$ for all parameters, showing data collapse onto a single curve, with the line representing the analytical formula in Eq. 14. (F) Scaled $Q$ vs. $\mathcal{A}_{s}$ for small and large noise strengths, showing a deviation from data collapse.
We examine temporal precision for varying parameters: $\mathcal{A}$, $\mu$, $k$, $D$, and $\gamma$, as shown in Fig. 2. The quality factor $Q$ varies non-monotonically with $\mathcal{A}$, peaking at an optimal $\mathcal{A}$, which is independent of $k$ but increases with $\mu$. The maximum $Q$ increases with larger $\mu$ and $k$, and $Q$ increases monotonically with both for a fixed $\mathcal{A}$. Although higher $\mathcal{A}$, $\mu$, and $k$ intuitively enhance precision, the decrease in $Q$ for large $\mathcal{A}$ is unexpected.
Oscillation quality deteriorates with increasing $D$ (Fig. 2(C)), but $Q$ remains non-monotonic with $\mathcal{A}$, and the optimal value appears independent of $D$. For thermal noise following the fluctuation-dissipation theorem, $Q$ does not depend on fluid viscosity $\eta$ (Fig. 2(D)), despite $T_{0}$ being $\eta$-dependent. However, for non-fluctuation-dissipation noise, precision decreases with $\eta$.
The data of $Q$ for different $\mu$, $k$, $\mathcal{A}$, $D$, and $\gamma$ can be collapsed on a single curve if $Q$ is scaled with $D\gamma/k\mu^{2}$ and plotted as a function of scaled amplitude $2\mathcal{A}/\mu$ (see Fig. 2 (E)). This implies the existence of the following scaling law:
$Q(\mathcal{A},\mu,k)=\frac{k\mu^{2}}{D\gamma}g\left(\frac{2\mathcal{A}}{\mu}\right).$ (7)
The scale factor $k\mu^{2}/D\gamma$ represents the ratio of two intrinsic energy scales: active energy $k\mu^{2}$ and effective temperature $D\gamma$. The quality factor $Q$ depends on the oscillation amplitude ${\cal A}$ through the scaling function $g$, which is a function of the dimensionless ratio ${\cal A}/\mu$. Other potential energy scales, such as $k{\cal A}^{2}$ or $k{\cal A}\mu$, do not influence the quality factor. Interestingly, for large noise strengths, the scaling behavior breaks down, as shown in Fig. 2(F). It suggests that such a scaling exists only for small noise, which is indeed true for axonemal beating.
@@
where $J$ corresponds to a current, here, which is given by the oscillation velocity. The above quantity is used in recent literature to establish a thermodynamic uncertainty relation, which sets an upper limit based on dissipation. The above definition of quality factor ${\cal Q}$ and $Q$ defined in Eq. 5 can differ by a proportionality constant, as shown in Fig. 3. The position-dependent switching between potentials introduces non-linearity, making direct analytical calculations difficult. To simplify, we focus on the Brownian motion within each potential and define the current as $J=\frac{2\mathcal{A}}{\mathcal{T}}$, where $\mathcal{T}$ is the time to travel between $-\mathcal{A}$ and $\mathcal{A}$ for $\sigma=1$ (or vice versa for $\sigma=-1$). Within SNA, this leads to:
@@
The relative variance in first-passage time (FPT) is given by $CV^{2}(\mathcal{T})=\frac{\langle\mathcal{T}^{2}\rangle-\langle\mathcal{T}\rangle^{2}}{\langle\mathcal{T}\rangle^{2}}$ [76, 77, 78, 79, 80], offering a clearer understanding of the system’s dynamics. FPT fluctuations, driven by switching events, introduce imprecision in the oscillations. By treating $\mathcal{T}$ as the first-passage time to reach $\mathcal{A}$ from $-\mathcal{A}$, we derive the formula for $CV^{2}(\mathcal{T})$ in the small fluctuation limit, as discussed in the following section.
Figure 3: The precision quality factors from two definitions are plotted against the scaled amplitude $\mathcal{A}_{s}$. Multiplying $\mathcal{Q}$ by 0.1 aligns the data, indicating the definitions differ only by a proportionality constant. Parameters: $\mu=2\,\mu m$, $k=4\,pN/\mu m$. The proportionality is more accurate at low noise (A) than at higher noise levels (B).
Figure 4: Relative variance of FPT, $CV^{2}(\mathcal{T})$: (A) $CV^{2}(\mathcal{T})$ vs. $\mathcal{A}$ for varying $\mu$. (B) $CV^{2}(\mathcal{T})$ vs. $\mathcal{A}$ for different $k$ values. (C) $CV^{2}(\mathcal{T})$ vs. $\mathcal{A}$ for varying $D$. (D) $CV^{2}(\mathcal{T})$ vs. $\mathcal{A}$ for different $\eta$. (E) Scaled $CV^{2}(\mathcal{T})$ vs. $\mathcal{A}_{s}=2\mathcal{A}/\mu$ for all parameters in (A)-(D), with the black line showing the analytical formula (Eq. 14). (F) Deviation from SNA results for higher noise strengths in the scaled plot. Parameters as in Fig. 2.
@@ III.3 First passage time (FPT) analysis
Due to the symmetry of the potentials, the FPT statistics for one potential are identical to those of the other. Let $\mathcal{T}$ denote the FPT, the time for the bead to reach $\mathcal{A}$ for the first time from $-\mathcal{A}$ in the harmonic potential $\frac{1}{2}kx(x-\mu)$. We aim to study the relative variance $CV^{2}$ in $\mathcal{T}$. The exact mathematical integral expressions are complex and limit analytical insight  [81, 77]. Using SNA, a method proven effective for FPT problems in gene expression [82, 83], we derive the relative variance (see Appendix A for details):
@@
$CV^{2}(\mathcal{T})\approx CV^{2}(x)\left(\frac{t}{\langle x\rangle}\frac{d\langle x\rangle}{dt}\right)^{-2}\Bigg{|}_{t=\langle\mathcal{T}\rangle},$ (10)
where $CV^{2}(x)=(\langle x^{2}\rangle-\langle x\rangle^{2})/\langle x\rangle^{2}$, is the relative variance in displacement at the mean FPT. Solving Eq. 2 for the first two moments in $x(t)$ (see Appendix A for a derivation), we get,
@@
$\langle x(t)\rangle{=}-\mathcal{A}\,\exp\left(-\frac{kt}{\gamma}\right){+}\frac{\mu}{2}\left[1-\exp\left(-\frac{kt}{\gamma}\right)\right],$ (11)
$\langle x^{2}(t)\rangle-\langle x(t)\rangle^{2}{=}\frac{D\gamma}{k}\left[1{-}\exp\left(-\frac{2kt}{\gamma}\right)\right].$ (12)
@@
Within mean-field ignoring stochasticity, one obtains the mean FPT $\langle\mathcal{T}\rangle=T_{0}/2$, half of the mean period. Using Eqs. 11 and  12 obtained at $t=\langle\mathcal{T}\rangle$ in Eq. 10, we find
$CV^{2}(\mathcal{T})=\frac{D\gamma}{k\mu^{2}}\frac{16\,\mathcal{A}_{s}}{\left[(1-\mathcal{A}_{s}^{2})\log\left(\frac{1+\mathcal{A}_{s}}{1-\mathcal{A}_{s}}\right)\right]^{2}},$ (13)
@@
where, $\mathcal{A}_{s}=2\mathcal{A}/\mu$, is the scaled amplitude. The expression above depicts a non-monotonic change in $CV^{2}(\mathcal{T})$ with $\mathcal{A}_{s}$, peaking at $\mathcal{A}_{s}=0.518$.
@@
The simulation results of $CV^{2}({\cal T})$ for various values of $k$, $\mu$, $D$, and $\gamma$ are shown in Fig. 4(A)-(D). All data collapse into a master curve when scaled by the factor $\frac{k\mu^{2}}{D\gamma}$ and plotted against ${\cal A}_{s}$, as described by Eq. 13, shown in Fig. 4(E). Since $Q\propto 1/CV^{2}({\cal T})$, this data collapse for $CV^{2}({\cal T})$ also explains the collapse observed for $Q$ in Fig. 2. The scaling function in Eq. (7) is given by
@@
up to a multiplicative constant. The plot of the scaling function in Fig. 2(E) shows good agreement with numerical results. However, at higher fluctuations, a departure from the data collapse is observed in both $Q$ and $CV^{2}({\cal T})$, as shown in Fig. 2(F) and Fig. 4(F).

### v2

@@ Abstract
Cilia and flagella are micron-sized filaments that actively beat with remarkable precision in a viscous medium, driving microorganism movement and efficient flow. We study the rower model to uncover how cilia activity and dissipation enable this precise motion. In this model, cilia motion is represented by a micro-bead’s Brownian movement between two distant harmonic potentials. At specific locations, energy pumps trigger potential switches, capturing cilia activity and generating oscillations. We quantify precision of oscillation using a quality factor, identifying its scaling with activity and oscillation amplitude, finding precision maximization at an optimal amplitude. The data collapse is not accurate for noisy oscillations. An exact analytic expression for the precision quality factor, based on first passage time fluctuations, and derived in the small noise approximation, explains its optimality and scaling. Energy budget analysis shows the quality factor’s consistency with the thermodynamic uncertainty relation. Finally, we demonstrate that asymmetric beating reduces oscillation precision compared to the symmetric model: although the optimal amplitude remains unchanged, the overall scaling of the quality factor depends explicitly on the asymmetry parameter.
@@ I Introduction
Motile cilia and flagella, hairlike appendages found on many unicellular organisms and epithelial tissues, actively beat to drive movement [33, 78, 10, 76]. Their structure consists of nine outer microtubule doublets and two central doublets [67], with sliding movement powered by dynein motors [73, 29]. This internal force causes microtubules to slide relative to each other [67, 59, 75, 14, 43, 85], generating complex oscillations in the surrounding fluid medium. These movements, often synchronized into metachronal waves, are essential for processes like mucociliary clearance [79, 78, 70] and the locomotion of microorganisms such as Paramecium, Opalina, and Volvox [80, 53, 12].
Cilia and flagella exhibit both coordinated movement in large groups and remarkable precision in their individual oscillations, despite fluctuations arising from the intrinsic stochastic activity of molecular motors and external noises from ambient fluid dynamics. The precision of these oscillations is quantified using a quality factor $Q$, a dimensionless ratio of coherence time to oscillation period [72, 5, 4, 74]. Values of $Q$ typically range from $\sim 30-100$ [65, 57, 30, 31, 52, 55, 77], indicating high precision. In Chlamydomonas reinhardtii, flagellar precision increases with the number of motor proteins. Experimental findings are summarized in a recent study [77].
@@
We investigate how the interplay between cilia activity and dissipation due to the surrounding viscous medium influences precision of oscillation, a topic of recent interest [55, 77]. To explore this, we first analyze the rower model numerically and then analytically in the symmetric case, quantifying precision using two related definitions of quality factor and linking it to fluctuations in the first passage time (FPT). We subsequently extend the analysis to encompass the asymmetric case. Our results show a clear data collapse and scaling of the quality factor, revealing a competition between activity and effective temperature. However, increasing noise strength causes deviations from the nice data collapse. Precision varies non-monotonically with oscillation amplitude, peaking at an optimal value. Applying a small noise approximation (SNA), we derive a closed-form expression for the FPT fluctuation and hence the quality factor, offering key insights into the role of activity and dissipation in precise beating. This expression explains the observed data collapse and scaling, and captures the precision peak at an optimal oscillation amplitude. As we show, the energy budget for achieving a desired precision, is consistent with the thermodynamic uncertainty relation.
@@ II The Rower Model
Here, $k$ is the stiffness, and $\mu$ is the separation between the minima of the two potentials, a natural length scale of the system. The motion is confined to the range $\pm\mathcal{A}$, where $2\mathcal{A}<\mu$. The bead undergoes Brownian motion only in the downhill regions of the potentials, and switching between the potentials at the terminal positions drives the bead in the opposite direction, creating sustained oscillations. Thermal fluctuations and other stochastic processes introduce noise into the oscillations. The over-damped motion of the bead is described by:
@@
where $\gamma=6\pi\eta a$ is the coefficient of drag force on the bead with radius $a$ due to the fluid with viscosity $\eta$, and $\xi(t)$ the white noise at time $t$. The noise satisfies $\langle\xi(t)\rangle=0$ and $\langle\xi(t_{1})\xi(t_{2})\rangle=2D\delta(t_{1}-t_{2})$, where $D$ denotes a translational diffusivity. Within this model, the diffusivity and drag coefficient can be used to define an effective temperature
@@
which, for active cilia, arises out of both thermal fluctuations and chemical processes involved in active energy pumping [66, 52, 86, 69]. We investigate temporal coherence in oscillations of this system. When the bead reaches a terminal position ($\pm\mathcal{A}$), the potential switches, injecting energy $\mu k\mathcal{A}$, which we refer to as the bead’s activity.
Figure 1: Oscillations in the rower model for cilia. (A) A schematic of the realistic beating of a cilium is depicted. (B) It illustrates the oscillation mechanism in the rower model. A micron-sized bead in a fluid medium moves in the downhills of two harmonic potentials (represented by $\sigma=\pm 1$) with stiffness $k$ and separated by a distance $\mu$. The motion is restricted between $\mathcal{A}$ to $\mathcal{A}$ with $2\mathcal{A}<\mu$. Once the bead reaches a terminal position ($\pm\mathcal{A}$), switching between the potential happens, implying the pumping of energy $\mu k\mathcal{A}$. (C) and (E) Typical oscillating trajectories for high noise ($D=0.02\mu m^{2}/s$) and low noise ($D=0.005\mu m^{2}/s$) strengths are shown. (D) and (F) Auto-correlation functions for trajectories for subplots (C) and (E) are shown. Other parameters used: $\mu=1\mu m$, $\mathcal{A}=0.25\mu m$, and $k=1.5pN/\mu m$.
Within a single potential well, the bead relaxes to equilibrium with a relaxation time $\tau_{d}=\gamma/k$, a natural time scale, in the absence of switching. The switching causes sustained oscillations of the cilium between $\mathcal{A}$ and $\mathcal{A}$ in this overdamped dynamics. The average period $T_{0}$ of the oscillations is obtained by solving the mean dynamics for each $\sigma$ separately, and is given by:
@@ III Results
In this section, we present definitions of quality factor describing the precision of oscillation and analyze the results, first numerically and then analytically within SNA.
@@ III.2 Precision quality factors of oscillations
The starting time $t_{0}$ should be large enough to avoid transients, and $\langle\cdot\rangle$ represents averaging over $t_{0}$ and trajectory ensembles. For noisy oscillations, $A(t)$ decays with lag time $t$ and can be fitted to $\cos(2\pi t/T)\exp(-t/\tau_{c})$, allowing estimation of $\tau_{c}$ and $T$, and calculation of $Q$. A higher $Q$ indicates better precision. In Fig. 1(C) and (E), typical trajectories for high ($D=0.02\ \mu m^{2}/s$) and low noise ($D=0.005\ \mu m^{2}/s$) are shown, with corresponding $A(t)$ in Fig. 1(D) and (F). For high $D$, $Q\approx 3$, and for low $D$, $Q\approx 12$.
The quality factor $Q$ generally depends on the parameters $\mathcal{A}$, $\mu$, $k$, $\gamma$, and $D$. However, simple dimensional analysis shows that $Q$ depends only on the scaled amplitude $\mathcal{A}_{s}=2\mathcal{A}/\mu$ and the dimensionless noise strength $1/\epsilon=D\gamma/(k\mu^{2})$. Using $\tau_{d}=\gamma/k$ as the characteristic time and $\mu/2$ as the length scale, the rower dynamics (Eq. 2) can be written in dimensionless form:
$\dfrac{dx^{\prime}}{dt^{\prime}}=-(x^{\prime}-\sigma)+\xi^{\prime}(t^{\prime})$
where $x^{\prime}=2x/\mu$ and $t^{\prime}=t/\tau_{d}$ and the white noise satisfies $\langle\xi^{\prime}(t_{1}^{\prime})\xi^{\prime}(t_{2}^{\prime})\rangle=(8/\epsilon)\,\delta(t_{1}^{\prime}-t_{2}^{\prime})$. The potential switches at $x^{\prime}=\pm\mathcal{A}_{s}$. While the average period $T$ depends only on $\mathcal{A}_{s}$ (Eq. 4), the coherence time depends on both $1/\epsilon$ and $\mathcal{A}_{s}$, so that $Q=Q(\mathcal{A}_{s},\epsilon)$. Its explicit functional form, however, is nontrivial.
@@
[…]
), (ii) $0.017$, $10.0$, $2.0$, and $6.0$ ($\triangle$), (iii) $0.017$, $10.0$, $2.0$, and $3.5$ ($\bigtriangledown$), (iv) $0.015$, $10.0$, $3.0$, and $2.0$ ($\square$), and (v) $0.01$, $7.4$, $2.0$, and 44 (
[…]
Numerical results reveal two key features of the quality factor $Q$. First, $Q$ varies non-monotonically with $\mathcal{A}_{s}$, peaking at an optimal value (Fig. 2). Second, for small noise, $Q$ data for different $\mu$, $k$, $\mathcal{A}$, $D$, and $\gamma$ collapse onto a single curve when scaled by $\epsilon$ and plotted against $\mathcal{A}_{s}$ (Fig. 2(A)), indicating the scaling law
$Q({\mathcal{A}_{s},\epsilon})=\epsilon g\left(\mathcal{A}_{s}\right).$ (7)
The existence of such a scaling is not obvious. While the general small-noise ansatz is $Q({\cal A}_{s},\epsilon)=\epsilon^{\psi}g({\cal A}_{s}/\epsilon^{\chi})$, our data show the simpler form with $\psi=1$ and $\chi=0$. For large noise, this scaling breaks down (Fig. 2(B)), with $g$ depending on both $\mathcal{A}_{s}$ and $\epsilon$. Thus, Eq. 7 holds only in the small-noise regime, relevant for axonemal beating.
@@
where $J$ corresponds to a current, here, which is given by the oscillation velocity. The above quantity is used in recent literature to establish a thermodynamic uncertainty relation, which sets an upper limit based on dissipation. The above definition of quality factor $\cal Q$} and $Q$ defined in Eq. 5 can differ by a proportionality constant, as shown in Fig. 3. The position-dependent switching between potentials introduces non-linearity, making direct analytical calculations difficult. To simplify, we focus on the Brownian motion within each potential and define the current as $J=\frac{2\mathcal{A}_{s}}{\mathcal{T}}$, where $\mathcal{T}$ is the (dimensionless) time to travel between $\mathcal{A}_{s}$ and $\mathcal{A}_{s}$ for $\sigma=1$ (or vice versa for $\sigma=-1$). Within SNA, this leads to:
@@
The relative variance in first-passage time (FPT) is given by $CV^{2}(\mathcal{T})=\frac{\langle\mathcal{T}^{2}\rangle-\langle\mathcal{T}\rangle^{2}}{\langle\mathcal{T}\rangle^{2}}$ [28, 34, 71, 51, 50], offering a clearer understanding of the system’s dynamics. FPT fluctuations, driven by switching events, introduce imprecision in the oscillations. By treating $\mathcal{T}$ as the first-passage time to reach $\mathcal{A}_{s}$ from $\mathcal{A}_{s}$, we derive the formula for $CV^{2}(\mathcal{T})$ in the small fluctuation limit, as discussed in the following section.
Figure 3: The precision quality factors from two definitions are plotted against the scaled amplitude $\mathcal{A}_{s}$. Multiplying $\mathcal{Q}$ by 0.1 aligns the data, indicating the definitions differ only by a proportionality constant. Parameters: $\mu=2\,\mu m$ $k=4\,pN/\mu m$. The proportionality is more accurate at low noise (A) than at higher noise levels (B).
@@ III.3 First passage time (FPT) analysis
Due to the symmetry of the potentials, the FPT statistics for one potential are identical to those of the other. Let $\mathcal{T}$ denote the FPT, the time (in dimensionless unit) for the bead to reach $\mathcal{A}_{s}$ for the first time from $\mathcal{A}_{s}$ in the harmonic potential $kx^{\prime}(x^{\prime}-1)/2$. We aim to study the relative variance $CV^{2}$ in $\mathcal{T}$. The exact expressions of the relative variance $CV^{2}(\mathcal{T})$ is given by  [51, 50]
$CV^{2}(\mathcal{T})$ $=\frac{\langle\mathcal{T}^{2}\rangle{-}\langle\mathcal{T}\rangle^{2}}{\langle\mathcal{T}\rangle^{2}}$
$=\frac{2\int_{y_{-}}^{\infty}{dy\,e^{y^{2}}[\operatorname{erfc(y)}]^{2}}\int_{y_{-}}^{y}{dz\,e^{z^{2}}\mathcal{H}(y_{+}{-}z)}}{\left[\int_{y_{-}}^{y_{+}}dy\;e^{y^{2}}\operatorname{erfc(y)}\right]^{2}},$
where $y_{\pm}=(1\pm\mathcal{A}_{s})/\sqrt{8/\epsilon}$, $\operatorname{erfc}(.)$ is the complementary error function, and $\mathcal{H}(.)$ is the Heaviside step function. As this expression involves nested integrals with error functions, analytical insight is limited, and the dependence on $\mathcal{A}_{s}$ and $\epsilon$ is nontrivial.
If the fluctuations in $x^{\prime}$ are small and the dynamics of the mean and variance are known, then an approximated formula for uncertainty in FPT can be obtained using the latter. Using SNA, a method proven effective for FPT problems in gene expression [9, 23], we derive the expression for the relative variance. Let us assume that the time $t^{\prime}$ for the bead to reach a position $x^{\prime}$ from an initial position fluctuates around the mean $\langle t^{\prime}\rangle$, i.e., $t^{\prime}=\langle t^{\prime}\rangle+\delta t^{\prime}$, where the fluctuation $\delta t^{\prime}$ is small. Then, ignoring higher-order terms in the Taylor expansion of $x^{\prime}(t^{\prime})$, we get
$x^{\prime}(t^{\prime})$ $\approx$ $\langle x^{\prime}(t^{\prime})\rangle+\delta t^{\prime}\frac{dx^{\prime}}{dt^{\prime}}\Big|_{t^{\prime}=\langle t^{\prime}\rangle}.$ (11)
Taking the square of both sides of Eq. 11 we obtain the expectation value
@@
$\langle\delta x^{\prime 2}\rangle$ $\approx$ $\langle\delta t^{\prime 2}\rangle\left(\frac{d\langle x^{\prime}\rangle}{dt^{\prime}}\right)^{2}\Big|_{t^{\prime}=\langle t^{\prime}\rangle},$ (12)
where, $\delta x^{\prime}=x^{\prime}(t^{\prime})-\langle x^{\prime}(t^{\prime})\rangle$. In terms of the squared coefficient of variation or relative variance, Eq. 12 can be written as
@@
$CV^{2}(\mathcal{T})\approx CV^{2}(x^{\prime})\left(\frac{t^{\prime}}{\langle x^{\prime}\rangle}\frac{d\langle x^{\prime}\rangle}{dt^{\prime}}\right)^{-2}\Bigg|_{t^{\prime}=\langle\mathcal{T}\rangle},$ (13)
where $CV^{2}(x^{\prime})=(\langle x^{\prime 2}\rangle-\langle x^{\prime}\rangle^{2})/\langle x^{\prime}\rangle^{2}$, is the relative variance in displacement at the mean FPT. Solving Eq. 2 for the first two moments in $x^{\prime}(t^{\prime})$, we get,
$\langle x^{\prime}(t^{\prime})\rangle{=}-\mathcal{A}_{s}\,\exp\left(-t^{\prime}\right){+}\left[1-\exp\left(-t^{\prime}\right)\right],$ (14)
$\langle x^{\prime 2}(t^{\prime})\rangle-\langle x^{\prime}(t^{\prime})\rangle^{2}{=}\frac{4}{\epsilon}\left[1{-}\exp\left(-2t^{\prime}\right)\right].$ (15)
@@
Within mean-field approximation, one obtains the mean FPT $\langle\mathcal{T}\rangle=T_{0}/(2\tau_{d})$, half of the mean period. Using Eqs. 14 and  15 obtained at $t^{\prime}=\langle\mathcal{T}\rangle$ in Eq. 13, we find
$CV^{2}(\mathcal{T})=\dfrac{1}{\epsilon}\frac{16\,\mathcal{A}_{s}}{\left[(1-\mathcal{A}_{s}^{2})\log\left(\frac{1+\mathcal{A}_{s}}{1-\mathcal{A}_{s}}\right)\right]^{2}}.$ (16)
@@
The expression above depicts a non-monotonic change in $CV^{2}(\mathcal{T})$ with $\mathcal{A}_{s}$, peaking at $\mathcal{A}_{s}=0.518$.
Figure 4: Relative variance of FPT, $CV^{2}(\mathcal{T})$: (A) Scaled $CV^{2}(\mathcal{T})$ vs. $\mathcal{A}_{s}=2\mathcal{A}/\mu$ for all parameters, with the black line showing the analytical formula (Eq. 17). (B) Deviation from SNA results for higher noise strengths in the scaled plot. The dashed lines are obtained solving the exact formula given by Eq. III.3. Parameters as in Fig. 2.
@@
For small noise strength, the simulation results of $CV^{2}({\cal T})$ for various values of $k$, $\mu$, $D$, and $\gamma$ collapse on to a master curve when scaled by the factor $\epsilon=\frac{k\mu^{2}}{D\gamma}$ and plotted against ${\cal A}_{s}$, as described by Eq. 16, shown in Fig. 4(A). Since $Q\propto 1/CV^{2}({\cal T})$, this data collapse for $CV^{2}({\cal T})$ also explains the collapse observed for $Q$ in Fig. 2. The scaling function in Eq. (7) is given by
@@
up to a multiplicative constant. The plot of the scaling function in Fig. 2 (A) shows good agreement with numerical results. However, at higher fluctuations, a departure from the data collapse is observed in both $Q$ and $CV^{2}({\cal T})$, as shown in Fig. 2(B) and Fig. 4(B). In Fig. 4(B), numerical results obtained from the exact expression, Eq. III.3, show excellent agreement with Langevin simulations across both noise regimes. At large $D$, both results deviate simultaneously from the low-noise (SNA) prediction, indicating the breakdown of the SNA and the associated data collapse at high noise strengths.